Monday 08/17/15

9:45 am

RNA library prep

setup

  • protocol from NEB
  • discuss plans with Jeff
  • request training on BioAnalyzer and/or TapeStation instrument for library QC
  • collecting kit components

RNA purification

  • run out 1 uL on a 1% agarose gel in TAE
  • treat samples with NEB DNase I (1 uL in 20 uL reaction, used RQ1 DNase buffer, don’t have an NEB one. validated the RQ1 is also a DNase I from Promega, guaranteed RNase free)
  • Purify samples on AMPure RNA XP beads.
  • performed all washes on column
  • gel results: (clearly degraded). RNA is a month old (stored at -80C, but used for several RTPCR runs.
    degraded_TotRNA
  • probably do better to start 1 clean fresh run with newly isolated samples.
  • we can run through the procedure today with these samples to get familiar with the protocol.

First strand synthesis

  • ran without actinomyocin (don’t have any). Already did DNA digestion so hopefully this isn’t necessary

Second strand synthesis

  • setup is easy
  • run bead cleanup using Aline beads from Jeff in place of AMPure XP beads

More steps

  • purify dsDNA (beads)
  • end prep
  • adapter ligation
  • purify DNA (beads)
    • this uses a lot of beads. Ordered more Aline beads from Aline biosciences
  • PCR enrichment
    • ran on QPCR machine with EvaGreen
    • 14 cycles, started to saturate, pulled off
    • testing sample by running out product on a gel
    • Typhoon isn’t working — images are perfectly 100% blank. no noise, no contrast, no background of gel.
  • still have 1 more bead clean up to do.

STORM analysis

  • compute z-calibration from surface beads for STORM2 for Ella
  • cancelled STORM imaging for tonight (insufficient time to setup a run before midnight, insufficient time to collect 2 color data if the staining works before 10am).
  • signed up for STORM4 for Thursday and Friday for live imaging with Steven. STORM4 might not work — the 100x objective went missing.

Ph KD immuno

  • finished staining
  • tested samples on Turnkey. WT shows decent staining. Ph in KD not detectable, but need nuclear counterstain to validate cell region.
  • maybe try imaging these on confocal tomorrow.

Editing

  • Editing recommendations for YF application
  • Editing Ph manuscript draft
  • got new Ph manuscript draft, will try to add to current one.
Posted in Summaries | Comments Off on Monday 08/17/15

Sunday 08/16/15

11:00 am – 7:45 pm

RNAi QPCR

  • Oligo cleanup purified T7 DNA
  • (probably could have done a standard DNA cleanup, I don’t need the supershort 30 bp stuff)
  • ran plate of QPCR
  • running out of primers.
    • Should order more primers for Pc, Ph, Antp, en. (done)
    • would be good to add some extra versions in as well.
    • still need new primers for ph-d at least
  • results look reasonable. Ph KD not at its highest
    QPCR_results_150816b QPCR_results_150816

previous data

  • this looks better with the raw DNA from yesterday, just I screwed up some of the primer labels
  • maybe can repeat again when I get new primers in next week.
    QPCR_results_150814b QPCR_results_150814
  • Ph KD efficiency = 0.87 +/- 0.02 (3 biological replicates, mean +/- SEM)

new RNAi

  • set up new RNAi for Ph + mock at large scale
  • plus culture maintenance
    • passage Kc cells in SFX
    • passage Kc cells in Sniedner’s
    • passage Psc-deletion cells
  • all cells will be used for RNA extraction and sequencing

RNAi Ph quantify structure effects

  • stain single-color control region for with new mock cells
  • stain new ANTC-P3 + G6-P1
  • stain new BXC-P3 + F6-P1
  • these P1 probes are both good and should be decent controls for staining
  • check staining on alternative BXC-P1 + F6-P3 this evening.
    • this one also failed
    • heat block is definetely mis-behaving.
    • might have screwed up this hybe too — temp dropped 2 degrees C when I added the slides

Ph KD quantification: New immuno staining

  • stained Ph-KD cells and mock cells with primary (1 hr RT)
  • rinsed 30 min
  • re-block (ready for rabbit secondaries).
Posted in Summaries | Comments Off on Sunday 08/16/15

Friday 08/14/15

9:00 am – 5:45 pm

Plan

  • need to do fewer review experiments at once so we can do them more carefully
  • should focus on the live imaging and Pc analysis
  • need to get good QPCR
  • need to set up deep-sequencing
  • need to analyze internal domains
  • need to do 2-color experiments (try L2 G05 G06).

Lab meeting

  • Jiang practice talk (see notes)

RNAi QPCR

  • started from RNA again, ran total RNA prep using oligoT primer
  • spiked in 2 uL RNase H as per the SS3 kit, incubated at 37
  • set up QPCR reactions
    • diluted DNA 1:5, then added 2.5 uL per tube.
  • revised quality controls on QPCR analysis script
  • most genes no amplification
    • Dfd amplified inall the Ph KD but none of the controls
  • probably not enough cDNA (or maybe inhibitor?)
    mostlyPoorAmplification

Meeting with XZ

  • live imaging plans:
    • move forward with analysis of SIM data
    • focus on the TIRF2-telo
  • ph
    • internal domain analysis
    • domain cross-boundary interaction
    • qPCR
    • sequencing
    • immuno-staining (?)
    • western (?)
  • other things are lower priority.

TIRF2 imaging

  • test imaging TIRF2-GFP fusion
  • these look very encouraging
  • clearly visible on Turnkey by eye
  • lots of spots, low nuclear background, clear signal.
    TIRFtelo1
Posted in Summaries | Comments Off on Friday 08/14/15

Protected: lab meeting 08/25/15

This content is password protected. To view it please enter your password below:

Posted in Lab Meeting | Comments Off on Protected: lab meeting 08/25/15

Thursday 08/13/15

9:30 am – 7:30 pm, 9:00 pm – 11:45 pm

Chromatin Project RNAi

RT reactions

  • protocol
  • 5 uL gene-specific primer mix
  • 10 uL total RNA, (diluted to 400 ng/uL)
  • 5 uL RT buffer
  • 1 uL dNTPs
  • 1 uL Maxima
  • 2 uL RNasin

master 10x

  • 50 uL RT buffer
  • 10 uL dNTPs
  • 10 uL Maxima
  • 20 uL RNasin

notes

  • noticed precipitate in my RT buffer. opened a new buffer. thawing now.
  • see same thing in new RT buffer. Didn’t notice before, but it’s pretty subtle
  • Primer mix: alpha-Tub84, Act5C, GADPH1, Pc, Ph-p, Antp, Abd-B, en
  • sample order:
    1. Ph 7-7 10 uL, –> 250 ng/uL
    2. mock 7-7 10 uL –> 250 ng/uL
    3. Ph 7-14 2.5 uL –> 210 ng/uL
    4. mock 7-14 2.5 uL –> 210 ng/uL
    5. Ph 8-8 10 uL –> 210 ng/uL
    6. mock 8-8 5 uL –> 450 ng/uL
    7. Pc 8-8 5 uL –> 250 ng/uL
    8. mock 6-29 2.5 uL –> 250 ng/uL

RT clean-up

  • try alkaline hydrolysis followed by oligo column-clean-up.
  • eluted in 20 uL. Spec’d all samples. conc’s above
  • dilute 5 into 20 uL ddH2O. use 2 uL per sample to set up.
  • try running with a no primer control

QPCR

Reaction set-up per well

  • 2 uL diluted DNA ~50 ng/uL
  • 1.25 uL EvaGreen
  • 12.5 uL Phusion 2x master mix
  • 2.5 uL primer combo (fwd/rev)
  • 6.75 uL ddH2O

Master mix (80x)

  • 100 uL EvaGreen
  • 1000 uL Phusion 2x master mix
  • 540 uL ddH2O
  • 20.5 uL per well

Plate setup

  • 8 DNA samples + 1 no template control
  • 8 gene-primer combos for each
  • 72 PCR reactions

Analysis

  • failed!! essentially indistinguishable similar amplification in all wells. We should have clearly orders of magnitude separated samples
  • probably accrued too much PCR product for these reactions in the pipettes etc.

Gel-troubleshooting

  • complete nonsense:
    TroubleshootingNonsense
  • stupid me. I mixed the foward and reverse primers already for the PCR. Can’t use both forward and reverse primers in the RT, this creates a mess.

To do still

  • upload review!
  • analyze SIM data

Other stuff done

  • Psc null cells, finished freezing down stock
  • send back shuttle form for EMBO -DONE
  • checked staining on BXC cell cycle — not that great via turnkey. Should analyze this data further
Posted in Summaries | Comments Off on Thursday 08/13/15

Wednesday 08/12/15

9:45 am – 11:30 pm

To do

  • submit review
  • book flights for EMBO (travel form / shuttle requests due on Friday)

Chromatin project

centromeric heterochromatin

  • downloaded dm3 heterchormatic assemblies from UCSC here: http://hgdownload-test.cse.ucsc.edu/goldenPath/dm3/chromosomes/
  • mapping previously imaged GREEN regions back to the chromosomes
    CentromericChromatinImaged_150812
  • download modENCODE chip data for heterchromatic regions
  • downloading data from modENCODE: http://data.modencode.org/?Organism=D.%20melanogaster
  • checking modENCODE data for agreement. Here is BXC for K79me3 (red), K27me3 (blue), K9me2 (green)
    modENCODE_Kc_ChIPdata
  • compare to van Steensel data we’ve been using:
    vanSteensel_Kc_ChIPdata_b
  • looks pretty good, let’s go on with this.
  • finished designing and assembling library. Will order tomorrow.

dual color labeling

  • 750-P3-BXC + 647-P1-L2F6 no spots :(
    • abbreviated wash. Try washing longer.
    • try G5 G6 instead

STORM

  • 750 laser desktop configuration file not running properly.

live imaging

  • construct 1: mmaple3-MCP in H293 cells
    • looks good! we see some telomeres
  • construct 2: mmaple3-PCP in H293 cells
    • also looks good, a bit more backgroundy
  • construct 3: PCP-tagRFP in H293
    • looks good. maybe best yet (though not STORMable)
    • took some movies (called tagRFP) can estimate positional stability.
  • construct 4: U2OS cells, gRNA co-transfect. TagRFP-PCP
    • a lot of uber bright nuclei with nuceleolus stain
    • not so good
Posted in Summaries | Comments Off on Wednesday 08/12/15

Monday 08/10/15

10:00 am – 5:15 pm, 8:30 pm – 1:15 am

tasks

  • pick up Ph antibody from Patrick at MGH (antibody gift from Jurg Muller)
  • ELYRA training. — with extreme laser power images looked promising. Otherwise images looked substantially less nice than the confocal
  • don’t have compatible version of Zen installed. Applied for download of latest version of Zen Light.

Cell cycle analysis

  • CycA looks promising (for G2).
    CycA_antibody
  • CycB stains (supposed to by cytoplasmic). G2/M marker. Destinction not so good. Stain generally dimmer as well.
    CycB_antibody
  • mPCNA (for s-phase) also looks promising
    mPCNA_stain

Data Analysis

  • processing Ph KD data
Posted in Summaries | Comments Off on Monday 08/10/15

Friday 08/07/15

10:00 am – 11:00 pm

No lab meeting

STORM

  • Fixed focus lock error with Hazen on STORM2, 10am – 12pm
  • Set up O/N imaging of L2F03F04
  • started analysis of STORM data

Analysis

  • analyzing live imaging data from last night
  • looks reasonable, still some misquantified cells. Should build a GUI to click through these

Chromatin project update meeting with Xiaowei

  • see google doc notes

Meeting notes

  • PRE density vs. length
  • plot PRE density relative to 1
  • not a substantial correlation
  • potentially a weak correlation
Posted in Summaries | Comments Off on Friday 08/07/15

Saturday 08/08/15

11:00 am – 1:15 am

PcG RNAi

New Knockdowns and cell culture

  • plate Ph, Pc, and mock RNAi cells onto coverglass
    • density looks okay modulo a decent number aren’t sticking
  • isolated cells for RNA extraction
  • fixed plated cells to coverglass for imaging
  • Psc null cells
    • starting to dye for not changing media (despite growing too slowly to be confluent).
    • replaced media on small plate and scraped plate (both these look pretty decent toward confluency).
    • combined media into new small vial, maybe detached cells will attach and settle (took > 24 hrs to attach last time)
    • made new Psc media (using modified Schneider’s from Lonzo and HI FBS).

Imaging BX-C in Ph-KD

  • finish imaging of L2F03F04
  • start imaging of L4E24

Analysis of BX-C internal domains in Ph-KD

  • still fiddling with export options for ChromatinCropper2
  • working on building ChromatinCropper2 GUI. This always takes longer than I think.
    • meanwhile I’m working out what I hope is a better general system for GUIs.

Cell cycle studies

  • prep cells for hybe
  • hybe BX-C to cells.
Posted in Summaries | Comments Off on Saturday 08/08/15

Thursday 08/06/15

10:00 am – 11:00 pm

To do

  • PRE correlation analysis
  • expression correlation analysis
  • design and Green domain probes
  • finish updates to ChromatinCropper2
  • analyze internal domain data

Correlation analyses

  • repeated PRE analysis as discussed:
    • recomputed using Cavalli/Tanay labs latest PRE predictions
    • examined distance from trend as function of PRE density
    • explored PRE density and number vs. volume (compared to length vs. volume)
    • could trivially extend to Rg
  • loaded expression data, see some weak correlation to explain outliers
  • also see (surprisingly) correlation with CTCF peak density to explain outliers
  • should combine these two.

STORM analysis

bead calibration

  • chromatic calibration still works okay with defaults
  • z-calibration: would like to rewrite this function. Less automatic selection of curves and more user input.

Live imaging

  • attempt repeat of live imaging
  • 2nd round looks good.
  • repeated issues with sample movement after pipetting
  • really should remember double-stick tape next time.
Posted in Summaries | Comments Off on Thursday 08/06/15